# Yunrui Li,a Hao Xu,a and Pengyu Honga

### v2

@@
[Uncaptioned image] Solvent-Aware 2D NMR Prediction via Multi-Tasking Pre-Training and Iterative Unsupervised Learning
@@
[Uncaptioned image] Nuclear magnetic resonance (NMR) spectroscopy is crucial across diverse scientific fields, revealing detailed structural information, electronic properties, and molecular dynamic insights. Accurate prediction of NMR peaks in a spectrum from molecular structures allows chemists to effectively evaluate candidate structures by comparing predictions with actual shifts in experimental NMR spectra. This process facilitates peak assignments, thereby aiding in verifying molecular structures or identifying discrepancies. Although significant progress has been made in predicting one-dimensional (1D) NMR with Machine Learning (ML) approaches, 2D NMR prediction remains a challenge due to the lack of an annotated 2D NMR training dataset. To address this gap, we propose an Iterative Unsupervised Learning (IUL) approach to train a machine learning model for predicting atomic 2D NMR cross peaks and annotating peaks in experimental 2D NMR spectra. Initially, the model undergoes a Multi-Task pre-Training (MTT) phase using a set of annotated 1D ${}^{1}\text{H}$ and ${}^{13}\text{C}$ NMR spectra. Then, the model is iteratively improved through a fine-tuning process with IUL, alternating between using the model to annotate the unlabeled 2D NMR data and refining the model using the newly generated annotations. Using the proposed approach, we trained our model on 19,000 Heteronuclear Single Quantum Coherence (HSQC) spectra, tested it on 500 HSQC spectra with expert annotations, and further compared it with two traditional methods (ChemDraw and Mestrenova) on another expert-annotated HSQC dataset. For HSQC cross peak prediction, our model achieves Mean Absolute Error (MAE) of 2.035 ppm and 0.163 ppm for ${}^{13}\text{C}$ shifts and ${}^{1}\text{H}$ shifts on the test dataset, respectively, and outperforms the conventional tools. This performance demonstrates not only the model’s capability in accurately predicting chemical shifts, but also its effectiveness in peak assignments for experimental HSQC spectra.
@@ 1 Introduction
In light of the aforementioned challenges and opportunities in interpreting HSQC spectra, we propose a novel solvent-aware machine learning model designed to predict HSQC cross peaks based on molecular structures, with the capability of peak assignment for experimental spectra, illustrated in Figure 1. Alongside a Graph Neural Network (GNN) module capturing structural nuances, the model incorporates a solvent encoder to effectively account for the impact of solvent environments on chemical shifts, which is essential for delivering accurate cross peak prediction and peak assignment of HSQC spectra. To tackle the lack of annotated HSQC data, we designed a two-step training process. First, the model is pre-trained on a labeled 1D NMR dataset via Multi-Task pre-Training (MTT), enabling it to learn a wide range of C–H interactions. Then, we implement the Iterative Unsupervised Learning (IUL) strategy that uses the unlabeled HSQC dataset to refine the model’s ability to accurately discern and label HSQC cross peaks. The model is thoroughly evaluated and compared to traditional tools like ChemDraw22 and Mestrenova23 using expert-annotated test datasets. It demonstrates superior accuracy across all molecular weight categories compared to these traditional tools.
Fig.  1: Illustration of model design (A) and training strategy ((B) and (C)). (A) The model takes a molecular structure and derives its atomic representations by a GNN. The solvent information is encoded into a latent representation via the Solvent encoder. The representation of each atom is concatenated with the solvent representation, which is then used to predict the atomic NMR shifts using an MLP component. The cross peak predictions are generated by linking the atomic predictions based on their connectivity in the molecular graph. (B) Model pertaining on the annotated 1D NMR dataset using MTT. (C) The pre-trained model is refined through the IUL process using the unlabeled HSQC dataset. The final output of the model has both the HSQC cross peaks and atom alignment.
@@ 2 Methods
In this section, we explain the components of our model and the training strategy in detail.
2.1 2D NMR prediction model
As illustrated in Figure 1(A), our model contains a GNN component for encoding molecular features and a solvent encoder component for embedding solvent information. The GNN component learns atomic embeddings that capture both the local and global chemical environments of each atom, which are essential for understanding the observed NMR chemical shifts. The learnt atom representations are expanded by the solvent embedding, and then are mapped to ${}^{13}\text{C}$ and ${}^{1}\text{H}$ cross peaks by a multilayer perceptron (MLP) component.
2.1.1 GNN
A molecule can be represented by a graph $G=(V,E)$, where $V$ is the node set representing atoms and $E$ is the edge set representing chemical bonds. Three features are provided for each node: atomic type, chirality, and hybridization. Also, two features are considered for each edge: bond type and bond direction. Bond types include Single, Double, Triple, and Aromatic, each reflecting a distinct configuration of electron sharing between atoms. Bond direction includes None, EndUpRight, and EndDownRight, primarily representing stereochemistry in double bonds. Each atom’s feature vector is embedded into a representation vector by a learnable encoder. Similarly, each edge’s feature vector is embedded into a representation vector of the same length by another learnable encoder. Then, a GNN model 24, 25, 26, 27, 28 utilizes the message passing mechanism to iteratively refine the representation of each node based on information from its neighbors and connected edges. This mechanism allows the learnt node representation to effectively capture structural context, reflecting the foundational principles of atomic interactions. Our implementation of the message passing mechanism is illustrated in 2. It iterates for a predefined number of layers $L$, facilitating the propagation of information throughout the graph. Consequently, each node can gradually accumulate information from a wider neighborhood across successive layers. This allow the final representation of each node to capture both local and global structural information. Our model features 5 GNN layers, with an atomic embedding dimension of 512.
@@ 2.1.1 GNN
2.1.2 Solvent encoder
Since the solvent has a profound impact on NMR chemical shifts, we incorporated a trainable solvent encoder component into our model to accurately capture this influence. We identified the following 9 principal solvent groups based on their prevalence in our dataset and domain-specific understandings of their distinct impacts on NMR shifts. These groups include trichloromethane, dimethyl sulfoxide, acetone, acids, benzene, methanol, pyridine, water, and an additional category to encompass any unspecified solvents from our dataset (termed "unknown"). The solvent encoder transforms each discrete solvent group $i$ into a unique, dense feature vector $S_{i}^{d}$, where $d$ is the embedding dimension. These learnable vectors are optimized alongside other model parameters during training, resulting in representations that accurately reflect the impact of each solvent class. Given the different sensitivities of carbon (C) and hydrogen (H) nuclei to solvent environments, we choose different embedding dimension $d$ to tailor the solvent effect modeling for each nuclei type. A larger embedding dimension $d$ allows the embedding to more effectively capturing the nuanced influence of solvents on NMR shifts. In our implementation, the solvent embedding dimension is 32.
2.1.3 Atomic NMR shift prediction
Finally, the embedding of each atom $h_{v}^{(L)}$ and the solvent embedding $S_{i}^{d}$ for each solvent class $i$ are concatenated to produce a holistic representation of the atom within the context of its molecular structure and the given solvent. This combined representation is subsequently processed by a MLP network to predict the NMR shifts for the atom:
$y_{v}=\text{MLP}(h_{v}^{(L)}\oplus S_{i}^{d})$ (1)
where $y_{v}$ is the predicted chemical shift of atom $v$, $h_{v}^{(L)}$ is the atom level embedding produced by GNN, $S^{d}$ is the solvent embedding, and $\oplus$ is the concatenation operation. By integrating solvent embedding and atomic embedding, the model effectively combines intrinsic molecular properties and solvent effects, enhancing its ability to predict atomic NMR shifts accurately.
Two separate MLP modules are used for predicting ${}^{13}\text{C}$ and ${}^{1}\text{H}$ shift in the cross peak predictions, respectively. Each C atom can bond up to 4 H atoms. When bonded to one, three, or four H atoms, a C atom typically shows only one cross peak in an experimental spectrum. However, when a C atom is connected to two H atoms, up to two cross peaks may be observed, depending on the chiral center. Consequently, a C atom can exhibit at most two ${}^{13}\text{C}$ and ${}^{1}\text{H}$ cross peaks. In light of this observation, one MLP module is dedicated to predicting the ${}^{13}\text{C}$ shifts and another MLP module for the corresponding ${}^{1}\text{H}$ shifts. For cross peak predictions, the ${}^{13}\text{C}$ shifts are predicted using the embeddings of C atoms. The corresponding ${}^{1}\text{H}$ shift predictions for each C atom incorporate aggregates of embeddings from all bonded H atoms, resulting in two predictions that are typically very similar when only one cross peak is theoretically possible. This design enhances the model’s accuracy in predicting ${}^{1}\text{H}$ shifts by leveraging the C atom-centered aggregation of the H atom context. By integrating the contextual dynamics around each C atom, the model provides a more detailed and accurate mapping of hydrogen environments, crucial for pinpointing precise cross peaks in complex HSQC spectra. In our implementation, we used 2 MLP layers, with the hidden dimensions to be 128 and 64 respectively.
2.2 Training strategy
The cross peaks are notably sparse in an HSQC spectrum, where typical resolutions for ${}^{13}\text{C}$ and ${}^{1}\text{H}$ shifts are 0.1 and 0.01 ppm, respectively. A typical HSQC spectrum can include  20,000 readings, covering ${}^{13}\text{C}$ shifts from 0 to 200 ppm and ${}^{1}\text{H}$ shifts from 0 to 10 ppm. However, almost all of these readings are zeros, with only a small fraction representing the potential cross peaks of C–H bonds, crucial for molecular structure analysis. Moreover, the scarcity of annotated HSQC data, particularly the labor-intensive annotations that link cross peaks to C–H bonds, makes model training difficult. To deal with this issue, we deployed MTT to pre-train the model using an extensive annotated 1D NMR dataset (Fig. 1B). This step acclimates the model with a broad range of molecular structures and their chemical shifts, and enables it to capture the intricate interplay between molecular structures and their NMR characteristics. Subsequently, we utilize the IUL strategy to refine the model further on the HSQC dataset (Fig. 1C). Through iterative cycles of prediction, annotation, and re-training, the model progressively enhances its understanding of the complex relationships and patterns within the HSQC spectra, thus improving its predictive accuracy and providing precise cross peak alignments. By combining the MTT and IUL, we extend our annotation capabilities from 1D to 2D data, thereby enhancing the model’s predictive power and utility as a robust tool for NMR spectra analysis.
2.2.1 Pre-training on 1D NMR data
In the pre-training phase, we utilized approximately 24,000 annotated 1D NMR data points. Among these, around 22,000 samples exclusively feature ${}^{13}\text{C}$ shifts, approximately 400 samples solely exhibit ${}^{1}\text{H}$ shifts, while roughly 1,600 samples contain both ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts. To train the model effectively for predicting both ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts, we adapt the MTT approach, which enables simultaneous training on multiple related tasks. When the input data contains ${}^{13}\text{C}$ shifts, the model predicts only carbon shifts and assesses the errors between the predicted and actual values. Conversely, when the data sample contains ${}^{1}\text{H}$ shifts, the hydrogen shift prediction module is activated. In both scenarios, the embeddings of ${}^{13}\text{C}$ and ${}^{1}\text{H}$ atoms in the GNN module are updated simultaneously, benefiting from the message passing mechanism. Therefore, the learnt representations implicitly contain a basic understanding of C–H relationships, essential for the interpretation of HSQC data. However, the relative scarcity of ${}^{1}\text{H}$ shift data, due to the difficulties in accurately obtaining and extracting peaks ${}^{1}\text{H}$ from experimental data, complicates the training process as focusing extensively on one type of shift could compromise the model’s ability to accurately predict the other. To handle this problem in the MTT training, we performed over-sampling on a subset of data that contain both ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts, and those containing only ${}^{1}\text{H}$ shifts. Consequently, the learned representations develop a fundamental understanding of C–H relationships, crucial for interpreting HSQC data effectively. This integration of learned atomic relationships streamlines the transition to HSQC cross peak predictions, thereby enhancing the model’s accuracy and efficiency in analyzing HSQC spectra.
2.2.2 Unsupervised fine-tuning on HSQC data
The model pre-trained on the 1D NMR dataset has limited ability to predict HSQC cross-peaks from molecular structures. The reason is two-fold. First, there is a substantial difference between the nature of 1D NMR data and HSQC data. The chemical shifts observed in HSQC cross peaks often do not correlate directly with their counterparts in 1D NMR spectra, particularly regarding ${}^{1}\text{H}$ shifts. For example, in the 1D NMR data, the chemical shifts of non-singlet peaks are averaged as ground truth, lacking the precision required by HSQC. Also, the variations in relaxation properties, coupling effects, and fluctuations in the local magnetic field can all contribute to the distinctions between 1D NMR and HSQC peaks. Moreover, experimental conditions and pulse sequences used in HSQC experiments can introduce slight deviations in chemical shift values compared to those observed in the ${}^{1}\text{H}$ spectrum. Second, the molecule distributions in our 1D NMR data and HSQC data exhibit significant differences. The HSQC dataset comprises 76.34% small molecules and 90.33% non-saccharides, whereas the 1D NMR dataset contains 98.80% small molecules and 99.95% non-saccharides.
Therefore, it is necessary to refine the pre-trained NMR prediction model on the HSQC dataset. However, the HSQC dataset is not annotated. In response, we implement an IUL training strategy (Fig. 1C), which iterates between (a) aligning cross peak prediction from the model with the experiment observations to annotate the HSQC data and (b) using the newly acquired annotations to fine-tune the NMR prediction model, until convergence.
2.2.3 Pseudo-annotation of HSQC
At the end of each round in the IUL process, the model’s predicted signals are aligned with the experimental observations to create pseudo-labels. In straightforward cases where the number of C–H bonds in a molecular graph matches the observed HSQC cross peaks, the Hungarian algorithm 29, 30 is used. This classic optimization technique solves assignment problems by minimizing the cost of matching a set of predictions to a set of observations. In the context of NMR analysis, the“cost” is defined as the discrepancy between the predicted chemical shifts and the actual shifts observed experimentally. By systematically reducing these differences, the Hungarian algorithm achieves an optimal one-to-one correspondence between predicted shift pairs and experimental signals, even in complex scenarios with potential signal overlap.
However, in most cases, the number of C–H bonds within a molecule exceeds the number of signals recorded, making peak alignment more difficult. This mismatch in numbers arises from several factors: firstly, rotational equivalence can reduce the number of signals, with a single peak representing all three C–H bonds for methyl groups; secondly, symmetrical molecular structures can result in a single detectable signal for multiple symmetric C–H bonds, as seen in benzene molecule where only one peak represents all six C–H bonds; lastly, in highly complex molecules, overlapping signals obscure some peaks, reducing the detectability of individual C–H bonds from experiments.
To overcome this issue, we utilize the graduated assignment algorithm 31, 15, which facilitates matching between graphs of different node counts, making it particularly suitable for this scenario. In this algorithm, our model’s predicted C–H shifts $(C^{i},H^{i})^{N}_{i=0}$ and the observed C–H signals $(C^{j},H^{j})^{M}_{j=0}$ of each molecule are conceptualized as points on a 2D plane, where $N$ and $M$ are numbers of predicted and observed C–H shifts respectively. These points are then treated as vertices in two fully connected graphs, $G_{1}$ for predicted shifts and $G_{2}$ for observed signals. The similarity between nodes is defined as the inverse of differences between predicted chemical shifts (node in $G_{1}$) and observed shifts (node in $G_{2}$). Specifically, for each predicted shift, we compute its difference with every observed shift, where a smaller difference indicates a higher similarity. The derive the assignment matrix $A$ where each element $A_{uv}\in\{0,1\}$ indicates whether node $u$ in $G_{1}$ matches with node $v$ in $G_{2}$, the algorithm first finds the soft matching matrix that relaxes the binary constraint $A_{uv}\in\{0,1\}$ to a continuous range $[0,1]$, then converts it into hard assignment in a greedy way, enabling one-to-many matching.
3 Results
@@ 3.1 Data
The pre-training dataset used in the MTT process is a 1D NMR dataset from NMRShiftDB28, which contains $\sim$24,000 annotated NMR spectra collected from 22,663 distinct molecules. The datasets used in the IUL process consist of a training dataset containing $\sim$19,000 experimental HSQC spectra and a validation dataset containing $\sim$5,000 HSQC spectra, collected from HMDB 32 and CH-NMR-NP 33. To quantitatively evaluate our model, we built a test dataset by randomly selecting 500 spectra and manually annotating them to establish the ground truth. These 500 spectra were randomly divided into 5 subsets of 100 spectra each, allowing us to conduct the assessment 5 times and evaluate the variation in performance. This test dataset was used to produce the results presented in Section 3.2, 3.4, 3.5, and 3.6. Additionally, to compare our model with two conventional tools (ChenDraw and Mestrenova) in chemistry, we created the second test dataset by randomly selecting a set of molecules and manually annotating their HSQC spectra to obtain 150 cross-peaks. The comparison results are presented in Section 3.3. The second test dataset is relatively small as it is labor-intensive to derive HSQC shifts from molecular formulas using these established tools. To ensure the coverage of the second test set, when selecting the molecules, we stratified all candidates into three weight groups (0-499 Dalton, 500-999 Dalton, and 1000+ Dalton) and randomly selected molecules from each weight group, resulting in 50 annotated cross peaks per groups.
3.2 Performance on HSQC cross peak prediction and assignment
Table 1 summarizes the performance of our model on the tasks of HSQC cross peak prediction and peak assignment using the 1st test dataset. Specifically, the model achieved an MAE of 2.06 ppm for ${}^{13}\text{C}$ shifts and 0.16 ppm for ${}^{1}\text{H}$ shifts. The robustness of our approach results in exceptionally high accuracy for peak assignments at both the molecular and peak levels, achieving 98.6% and 99.79% respectively. Our model is said to be correct on a molecule if all cross peaks of the molecule are correctly assigned. The peak level accuracy is calculated as the ratio of correctly assigned cross peaks to the total number of cross peaks. Figure 3 illustrates an example of the model outcome of peak prediction and assignment, using inputs of molecule and solvent.
Table 1: Performance on HSQC cross peak prediction and peak assignment for the 1st test set. This table reports the Mean Absolute Error (MAE) and the accuracy of peak assignments (the pseudo annotations) produced by our approach. The numbers in the parentheses are the standard deviations.
MAE (ppm) Peak Assignment Accuracy (%)
${}^{13}\text{C}$ Shift ${}^{1}\text{H}$ Shift Molecule Peaks
2.035 0.163 98.6 99.79
(0.088) (0.003) (0.55) (0.12)
Fig.  3: An example of using our model to accurately predict cross peaks and align them with experimental signals. The molecule is shown at the top-left, where each C–H bond is labeled with a numerical identifier. Notably, the symmetric pairs of bonds (labeled as “2”, “3”, and “4”) are each expected to generate a single HSQC cross peak due to their structural equivalence. The HSQC cross peaks predicted by our model (in orange) and their alignments to the experimental observations (in blue) are plotted in the right. The alignments are indicated by the dash circles.
3.3 Comparison with traditional tools
In organic chemistry, simulating HSQC spectra is crucial for analyzing experimental HSQC spectra, as it assists researchers in assigning the observed cross peaks to the C–H bonds in target molecules. Traditional approaches, including software solutions such as ChemDraw22, and Mestrenova23, have long served as the primary resources for this task. We compared our model with ChemDraw and Mestrenova on the second test dataset. The results (see Table 2) clearly demonstrate the superiority of our model. Figure 4 illustrates two examples with different molecular sizes, showing the advantage of our model.
Molecular ML Model MAE ChemDraw MAE Mestrenova MAE
Weight ${}^{13}\text{C}$ ${}^{1}\text{H}$ ${}^{13}\text{C}$ ${}^{1}\text{H}$ ${}^{13}\text{C}$ ${}^{1}\text{H}$
0 - 499 1.128 0.095 1.391 0.292 1.293 0.141
500 - 999 1.031 0.091 2.362 0.215 1.699 0.112
1000+ 3.891 0.171 6.862 0.601 7.748 0.409
Overall 2.003 0.119 3.529 0.367 3.559 0.219
Table 2: Performance comparison between our proposed model and established traditional tools on the second test dataset. Our model performs better across all molecular weight categories.
Fig.  4: Comparing our model, ChemDraw, and Mesteronva on two typical examples. A small molecule (a) with weight of $\sim$250 Dalton and a larger molecule (b) with weight of $\sim$500 Dalton. The observed experimental signals and the predicted signals are colored in blue and orange, respectively. The prediction error (MAE) is shown in the bottom right corner of each plot. Our model performs better than ChemDraw and Mesteronva, and particularly excels in handling large molecules with complex conformations.
3.4 Performance by molecular weights
Molecular weight is a general indicator of a molecule’s complexity, encompassing varied geometries, bonding patterns, and the presence of isomers. These factors contribute to increased intramolecular interactions, resulting in spectral complexity such as closely spaced peaks or overlapping signals. Additionally, the increased number of spin-spin interactions within larger molecules necessitates more advanced NMR techniques to achieve sufficient resolution.34 Furthermore, solubility issues can lead to weak signals, further complicating spectral analysis. Consequently, interpreting HSQC spectra for medium and large molecules is challenging. Therefore, there is a pressing need for a model that can effectively predict and analyze HSQC spectra for these complex molecules.
To address this challenge, we evaluated the robustness of our model on 1st test dataset comprising complex medium and large molecules, testing its performance across a range of molecular weights. The results are stratified by molecular weights (MW) in Figure 5. The test molecules are grouped into three categories: small (MW < 500 daltons), medium (500 <= MW < 1000 daltons), and large (1000 daltons <= MW). On the task of predicting ${}^{1}\text{H}$ shifts of HSQC cross peaks, the model performs comparably across all groups, achieving excellent MAEs of 0.16 - 0.17 ppm. On the task of predicting ${}^{13}\text{C}$ shifts, the model achieves an MAE of 1.91 ppm for medium-sized molecules. Our model demonstrates a good generalization power on large molecules and achieves a MAE of 2.14ppm on predicting ${}^{13}\text{C}$ shifts for this category, despite the training data containing only a small proportion of large molecules ($\sim$2%).

### v3

@@
[Uncaptioned image] AI-enabled prediction of NMR spectroscopy: Deducing 2-D NMR of carbohydrate
@@
[Uncaptioned image] In the dynamic field of nuclear magnetic resonance (NMR) spectroscopy, artificial intelligence (AI) has ushered in a transformative era for molecular studies. AI-driven NMR prediction, powered by advanced machine learning and predictive algorithms, has fundamentally reshaped the interpretation of NMR spectra. This innovation empowers us to forecast spectral patterns swiftly and accurately across a broad spectrum of molecular structures. Furthermore, the advent of generative modeling offers a groundbreaking approach, making it feasible to make informed prediction of 2D NMR from chemical language (such as SMILES, IUPAC Name). Our method mirrors the multifaceted nature of NMR imaging experiments, producing 2D NMRs for the same molecule based on different conditions, such as solvents and temperatures. Our methodology is versatile, catering to both monosaccharide-derived small molecules, oligosaccharides and large polysaccharides. A deeper exploration of the discrepancies in these predictions can provide insights into the influence of elements such as functional groups, repeating units, and the modification of the monomers on the outcomes. Given the complex nature involved in the generation of 2D NMRs, our objective is to fully leverage the potential of AI to enhance the precision, efficiency, and comprehensibility of NMR spectral analysis, ultimately advancing both the field of NMR spectroscopy and the broader realm of molecular research.
@@ 1 Introduction
In light of the aforementioned challenges and opportunities in interpreting HSQC spectra, we propose a novel solvent-aware machine learning model designed to predict HSQC cross peaks based on molecular structures, with the capability of peak assignment for experimental spectra, illustrated in Figure 1. Alongside a Graph Neural Network (GNN) module capturing structural nuances, the model incorporates a solvent encoder to effectively account for the impact of solvent environments on chemical shifts, which is essential for delivering accurate cross peak prediction and peak assignment of HSQC spectra. To tackle the lack of annotated HSQC data, we designed a two-step training process. First, the model is pre-trained on a labeled 1D NMR dataset, enabling it to learn a wide range of C–H interactions. Then, we implement an unsupervised strategy that uses the unlabeled HSQC dataset to refine the model’s ability to accurately discern and label HSQC cross peaks. The model is thoroughly evaluated and compared to traditional tools like ChemDraw22 and Mestrenova23 using expert-annotated test datasets. It demonstrates superior accuracy across all molecular weight categories compared to these traditional tools.
Fig.  1: Illustration of HSQC prediction and alignment pipeline. The model takes a molecular structure and derives its atomic representations. The final output of the model has both the HSQC cross peaks and atom alignment.
@@ 2 Methods
Our model integrates several components to facilitate the understanding of NMR data through innovative computational approaches. At its core, the model employs a Graph Neural Network (GNN) for encoding molecular features alongside a solvent encoder for capturing the effects of different solvents on NMR chemical shifts. The combination of these components allows the model to generate detailed predictions about molecular behavior in various environments, crucial for interpreting complex NMR data.
The GNN, as illustrated in Figure 2 is designed to learn detailed embeddings of atoms within a molecule, capturing both local and broader chemical environments which are critical for predicting NMR chemical shifts. This is achieved by representing molecules as graphs where nodes are atoms and edges represent bonds, with various attributes such as atomic type and bond characteristics enhancing the model’s understanding of molecular structure.
The model’s ability to integrate detailed molecular structure with solvent effects makes it a powerful tool for analyzing NMR data. This holistic approach is crucial for advancing our understanding of molecular interactions in different chemical environments, which is key for numerous applications in chemistry and related fields.
@@
3 Results and discussion
@@ 3.1 Data
Two datasets are used in the training process. The first one is the NMRShiftDB28 dataset which contains $\sim$24,000 annotated NMR spectra collected from 22,663 distinct molecules. The second dataset is collected from HMDB 24 and CH-NMR-NP 25, divided into a training dataset containing $\sim$19,000 experimental HSQC spectra and a validation dataset containing $\sim$5,000 HSQC spectra.
3.2 Pre-training performance
For 1D NMR data, our model simultaneously predict ${}^{1}\text{H}$ and ${}^{13}\text{C}$ NMR spectra using a single model framework. Our training and validation data includes shifts for ${}^{13}\text{C}$, ${}^{1}\text{H}$, or both, with the majority of the samples containing only ${}^{13}\text{C}$, a smaller subset of samples with both ${}^{1}\text{H}$ and ${}^{13}\text{C}$ shifts and only a few samples containing ${}^{1}\text{H}$ shifts. Despite the small portion of the ${}^{1}\text{H}$ shifts in the dataset, our model achieved impressive mean absolute errors (MAEs) of 0.210 ppm for ${}^{1}\text{H}$ NMR prediction and 2.228 ppm for ${}^{13}\text{C}$ NMR prediction, demonstrating the effectiveness of our approach. The success of our model can be attributed to GNN and message passing, which allows the model to learn ${}^{1}\text{H}$ atom features not only from directly prediction the ${}^{1}\text{H}$ shifts, but also from the connecting ${}^{13}\text{C}$ features through message-passing mechanisms. This approach helps the model to implicitly learn the C–H interactions which laid a solid foundation for the HSQC training in the next step. Using this method, our framework surpasses the data limitation from experiments where ${}^{13}\text{C}$ signals are easier to obtain than ${}^{1}\text{H}$, or vice versa. It offers a flexible procedure to utilize experimental data and learn molecular dynamics from them as much as possible. Overall, these results highlight the potential of multitasking training in enhancing the accuracy and performance of NMR prediction models.
Our 1D NMR data training incorporates two specifically designed features to optimize the model for 2D NMR predictions, albeit without immediate benefits for 1D NMR outcomes. Firstly, in our dataset for 1D NMR, solvent labels are scarcely provided. Consequently, we handle the solvent class as “unknown” during training. This approach yields no observable enhancement in performance compared to a model that does not account for solvent information. However, it facilitates a seamless adaptation to the HSQC dataset, characterized by its varied solvent labels, preparing the model for a more nuanced solvent-aware prediction in 2D NMR scenarios. Additionally, due to the differing distribution of the HSQC dataset (76.34% small molecules and 90.33% non-saccharides) compared to the 1D NMR datasets (98.80% small molecules and 99.95% non-saccharides), we anticipate challenges in transferring learning or fine-tuning the model for HSQC prediction. To address this, we intend to use fewer epochs, specifically 200 epochs, as opposed to the 500 epochs used in MPNN work 17. To reduce model overfitting on 1D data, we also left 20% data as validation and only 80% for training, compared to 5% and 95% data split in the MPNN work. This modification, while possibly resulting in slightly elevated MAEs for 1D predictions, enhances the model’s flexibility for HSQC fine-tuning and aligns with our strategic objective of creating a solvent-aware model suited for HSQC spectrum analysis. This balance between predictive accuracy and computational expediency is pivotal, enhancing our model’s overall feasibility and effectiveness for intricate HSQC predictions. Table 1 shows the model’s performance compared to benchmarak methods on 1D NMR dataset.
Table 1: MAEs for representative methods of 1D NMR prediction
Method 1H MAEs (ppm) 13C MAEs(ppm)
HOSE 8 0.29 3.52
DFT 11 0.23 2.9
MPNN 17 0.224 1.355
GNN + solvent 0.210 2.228
3.3 Convergence during self-training
Self-training is an iterative process aimed at minimizing validation errors while avoiding overfitting. We conducted a few iterations, as illustrated in Figure 3. Across five iterations, we observed a significant decrease in validation errors, particularly in the initial stages. In the first iteration, MAEs of C-shifts decreased by approximately 1.1 ppm, and MAEs of H-shifts decreased by about 0.12 ppm. Subsequently, in the second iteration, there was a decrease in MAEs of C-shifts by roughly 0.3 ppm and MAEs of H-shifts by around 0.1 ppm. After 5 iterations, the improvement becomes marginal between iterations. Consequently, we conclude that the process has converged. The MAE of model prediction error is shown in Table 2. Upon the convergence, the model is capable of producing accurate HSQC predictions while concurrently generating the atom alignment for each C–H bond. In Figure 4, we demonstrate an example of HSQC prediction and peak assignments.
Fig.  3: Error curve for ${}^{13}\text{C}$ and ${}^{1}\text{H}$ during iterative self-training phases. The pre-trained error shows direct application of the pre-trained model on 1D data to predict HSQC shifts. After 4 rounds of iterative self-training, the error curve gradually flattens, indicating the model has reaches satisfactory performance.
Table 2: MAEs of 2D NMR prediction after convergence
Dimension Training MAEs (ppm) Validation MAEs (ppm)
C-shifts 1.638 1.997
H-shifts 0.131 0.161
Fig.  4: Example of model usage on HSQC prediction and peak assignment. The molecule is shown in the left, where each C–H bond is labelled. The model’s HSQC spectrum prediction is shown in the right, with the model assigning accurate labels to these predictions. Notably, the model adeptly recognizes and accounts for symmetry within the molecule, assigning identical C–H shifts to symmetrical bond sets. This exemplifies the model’s advanced capability to discern and reflect molecular conformation in its predictions accurately.
3.4 Post-analysis of model performance
In this section, we highlight the model’s effectiveness across different molecular categories, emphasizing its performance on medium to large molecules and saccharides.
3.4.1 Performance regarding molecular weight
The interpretation of HSQC spectra for medium to large molecules often presents significant challenges, as their considerable size introduces greater structural complexity and various isomeric configurations. This complexity introduce more intramolecular interactions, and often results in crowded NMR spectra with overlapping signals, making it difficult to resolve individual peaks. Additionally, the increased number of spin-spin interactions within large molecules demands more advanced NMR techniques to achieve sufficient resolution. The longer relaxation times associated with larger molecules slow down experimental procedures, requiring more sensitive NMR equipment. Furthermore, solubility issues can lead to weak signals, complicating spectral analysis even further. Therefore, there is a pressing need for a model proficient in predicting and analyzing HSQC spectra for medium to large molecules.
As demonstrated in Figure 5, our model achieves significant accuracy in predicting shifts for both medium and large molecules. We categorize the dataset into three segments based on molecular weight (MW), utilizing 500 and 1000 daltons as the dividing thresholds. For small molecules (MW < 500 daltons), the HSQC prediction yields a MAE of 1.75 for ${}^{13}\text{C}$ shifts and 0.14 for ${}^{1}\text{H}$ shifts. For medium molecules (500 < MW < 1000 daltons), the HSQC prediction results in an MAE of 1.81 for ${}^{13}\text{C}$ shifts and 0.16 for ${}^{1}\text{H}$ shifts. For large molecules (MW > 1000 daltons), the HSQC prediction shows an MAE of 2.18 for ${}^{13}\text{C}$ shifts and 0.18 for ${}^{1}\text{H}$ shifts. The predictions for small and medium molecules exhibit comparable accuracy, with a slight reduction in precision for larger molecules, likely due to their rarity in the training dataset. Notably, large molecules constitute merely 2% of our dataset, reflecting the scarcity and difficulty in obtaining high-quality 2D NMR data for such molecules. Overall, our model demonstrates stable performance across different sizes of molecules, underscoring the effectiveness and promise of our framework in delivering precise and reliable 2D NMR shift predictions.